# Word-Level Emotional Expression Control in Zero-Shot Text-to-Speech Synthesis

### v1

@@ 1 Introduction
Humans possess the ability to regulate emotional expression during speech flexibly scherer1995expression . To simulate this expressive capability, recent advances in text-to-speech synthesis (TTS) have increasingly focused on controllable generation of various aspects of speech, such as timbre, emotion, and speaking rate xie2024towards . Such control is a key objective in the development of human-like and expressive TTS.
Most current TTS models exhibit zero-shot capabilities, enabling them to synthesize speech from text while cloning attributes such as timbre, emotion, and speaking rate from a reference speech sample kharitonov2023speak ; wang2023neural ; ju2024naturalspeech . Despite these advances, as shown in Figure 1, emotional and speaking rate control in current models is typically limited to the utterance level.
This differs significantly from how humans naturally express emotion in speech. Unlike global speaker identity, emotional expression and speaking rate are dynamic and often vary within a single sentence mozziconacci2002prosody ; guan2018speech . Therefore, word-level control of these factors is essential for achieving more natural and expressive speech synthesis 9746323 . To address this limitation, some approaches have proposed phoneme-level emotion prediction from target text to guide expressive synthesis tang2024ed ; 9383524 ; du21b_interspeech . While these methods show potential for word-level emotion control, relying solely on text makes it difficult to capture essential acoustic cues such as prosody and intensity, which are vital to emotional expression control chen2022fine ; chandra2024exploring ; zhao2023emotion . To address this limitation, recent studies such as ELaTE kanda2024making and EmoCtrl-TTS wu2024laugh have demonstrated that reference speech with emotional content can support intra-utterance control of time-varying expressive patterns, such as transitions from laughter to crying. These works reflect a growing interest in TTS with word-level control over both emotion and speaking rate, but they also underscore several fundamental challenges. First, word-level expression control requires multiple emotional speech prompts, which introduces the challenge of guiding the model to attend to the appropriate emotion at each word. In addition, current methods for fine-grained expression control rely on large-scale emotional speech datasets with time-aligned emotion transitions. However, such datasets are limited in both scale and accessibility rathi2024analyzing , making fine-grained control even more difficult to realize in practice. These challenges lead us to ask: Is it possible to achieve effective word-level control of both emotion and speaking rate without relying on speech datasets containing emotion or speed transitions?
@@
We conduct comprehensive experiments to validate the effectiveness of our proposed framework. Results show that our method enables a pretrained zero-shot TTS model to achieve SOTA performance in word-level emotional expression control, while preserving its original zero-shot capabilities. Ablation studies further confirm the contribution of each key design component. Our samples are available at https://anonymousdemo999.github.io/.
@@ 2 Related Work
The development of controllable TTS, particularly for emotional expression control, depends heavily on high-quality emotional speech datasets zhu2024metts_emodata ; guo2023emodiff_emodata ; yang2025emovoicellmbasedemotionaltexttospeech_emodata . While public corpora such as ESD zhou2021seen , IEMOCAP busso2008iemocap , and CREMA-D cao2014crema are available, they primarily provide utterance-level annotations and lack word-level or time-aligned emotional labels. These datasets are also limited in size and diversity, often consisting of scripted speech and covering a narrow range of emotions and speakers ma24b_interspeech . More importantly, emotional datasets with intra-sentence variation, which are essential for learning word-level control, remain extremely scarce and are typically restricted to private use kanda2024making . Creating such datasets is expensive, requiring detailed word- or frame-level annotation and subjective emotional labeling liu2024emphasisrenderingconversationaltexttospeech . This lack of fine-grained emotional data poses a major challenge for training models capable of word-level expressive TTS.
@@
Most controllable TTS systems support only utterance-level control, where a single label or reference speech governs the entire sentence anastassiou2024seed ; du2024cosyvoice . To achieve word-level control, some methods attempt to predict frame- or phoneme-level emotional indicators from text alone phone_level_tts0 ; phone_level_tts1 ; phone_level_tts2 , but they often fail to capture expressive variability due to the lack of acoustic cues such as intensity and prosody tang2024ed ; 9383524 ; du21b_interspeech . Other approaches, such as ELaTE kanda2024making and EmoCtrl-TTS wu2024laugh , introduce emotional reference speech to enable intra-utterance control of specific expressive patterns like laughter or crying. While these represent progress, they are typically limited in expressiveness or rely on large-scale emotional datasets that are rarely publicly available. Consequently, achieving general and flexible word-level control over both emotion and speaking rate remains a major challenge.
@@
Self-training has become a promising approach for low-resource speech signal processing, enabling knowledge transfer without fine-grained datasets zoph2020rethinking ; amini2025self . While it has been applied to tasks like speaker adaptation khurana2021unsupervised , paralinguistic modeling yang2024frame , and speech translation pino20_interspeech ; fang-etal-2022-stemm , its use for fine-grained emotional control in TTS remains unexplored, especially without detailed expressive labels. To address the scarcity of fine-grained datasets for word-level expressive control, we propose a self-training framework where a teacher model with multi-round inference, transition smoothing, and dynamic speed control generates expressive pseudo-labels. A student model, sharing the teacher’s backbone, is then fine-tuned under its supervision to perform word-level emotion and speaking rate control through a simplified end-to-end inference process, using only a small public dataset without intra-sentence emotion or speed transitions.
@@ 3.2.1 Word-Level Emotion Control
As discussed in Section 2, current TTS models can perform utterance-level emotion and speaker cloning. Building on this, we adopt the high-performance CosyVoice2 du2024cosyvoice2 as our backbone (details of the backbone architecture are provided in Appendix A) and propose a multi-round inference strategy, where the model synthesizes multiple segments using different emotional prompts to achieve word-level emotion control. While this approach enables flexible emotional modulation, it often causes unnatural acoustic discontinuities at segment boundaries. To address this, we introduce a transition-smoothing mechanism that improves coherence across inference rounds, as illustrated in Figure 3. Without modifying CosyVoice2, we append a lightweight content aligner, composed of non-causal Transformer vaswani2017attention and convolutional layers. Trained on ASR data, this module predicts the corresponding text token for each speech token and requires no emotional supervision. During inference, the input text is segmented based on a user-defined emotion plan. At each inference round, the final text and speech tokens from the previous round are appended to the current prompt, forming an explicit tail-to-head linkage. This aligns naturally with CosyVoice2’s continuation-style generation borsos2023audiolm ; valle , enabling smooth and coherent emotional transitions.
@@ 3.3 Self-Training
In the previous section, we enabled word-level control of emotion and speaking rate by introducing a multi-round inference framework for CosyVoice2 du2024cosyvoice2 . However, components such as the non-causal content aligner, multi-round inference, and tail-to-head linkage introduce significant inference complexity. To reduce this overhead while preserving controllability, we adopt a self-training strategy. As shown in Figure 4, the enhanced first-stage model serves as a teacher to supervise the original TTS model. The student model, equipped with a dynamic emotional attention bias, learns to achieve word-level emotion and speaking rate control through a simplified end-to-end inference.
@@ 3.3.1 Self-Training with Teacher-Generated Emotion-Transition Speech
Our teacher model achieves word-level control of emotion and speaking rate without modifying the original TTS parameters, relying instead on a complex inference pipeline with dynamic speed control and multi-round generation. To transfer this fine-grained control ability to a simplified end-to-end model, we propose a self-training strategy. Specifically, the 1st-stage teacher model guides the student model to learn word-level controllability. We first use GPT-4o hurst2024gpt to generate emotion-transition text sequences (details are shown in Appendix C), which are paired with public emotional speech samples (without emotion transitions) as prompts. The teacher then synthesizes speech with word-level variation in emotion and speaking rate. These outputs are filtered based on character accuracy and expressive similarity (details are introduced in Appendix D), and the student model is fine-tuned on the filtered supervisions with a small learning rate. This enables word-level emotional expression control during inference without requiring multi-round generation or dynamic concatenation.
@@ 3.3.2 Dynamic Emotional Attention Bias
While the above data formatting preserves CosyVoice2’s generalization by avoiding interference with learned knowledge, it introduces a new challenge: during synthesis, the model may incorrectly attend to emotion-inconsistent prompts. For instance, when generating speech aligned with Emotion I, attention may drift toward prompts labeled with Emotion II, leading to emotional inconsistency and degraded synthesis quality. To address this, we propose a dynamic attention bias mechanism that constrains the model’s focus to emotion-relevant prompt regions based on the predicted emotional trajectory. Concretely, we introduce a causal lightweight Transformer to predict token-level emotion labels for each speech token from historical context. Using the predicted emotion sequence, we introduce a dynamic attention bias mechanism at each Transformer layer. We first concatenate the current text-speech representation with the predicted emotion features and project it through a linear layer. The output is processed in two ways: one path adds a residual and feeds into the next layer, while the other is passed to an MLP popescu2009multilayer and softmax to produce a weight vector . The is then used to compute a dynamic attention bias by linearly combining seven predefined attention bias templates (see Appendix E for details). The resulting bias is computed as:
@@ 3.4 Detail Training Setup
We use forced alignment mcauliffe2017montreal to generate token-level alignments between transcripts and speech, which serve as supervision for the content aligner. The TTS model remains frozen throughout this stage. Training of the content aligner is conducted without multi-round forwards. Let and denote the input text and speech token sequences, denote the aligned target token sequence, where each label corresponds to one of token classes. Let be the binary label sequence for content boundary detection. The content aligner is jointly trained with a token-level content classification loss and a binary boundary detection loss:
where and denote the last frame indices for text and speech, and is the total number of speech tokens. The learnable parameters correspond to the content aligner, while is the frozen TTS model parameter used during forward propagation. We also apply class weighting during loss computation to reduce the impact of overrepresented silence tokens and address the imbalance in boundary label distribution he2009learning .
@@ 4.1 Experimental Setup
In the first stage, the content aligner is trained on 200 hours of non-emotional English-Chinese speech from LibriSpeech-100-Clean panayotov2015librispeech and AISHELL-1 bu2017aishell . In the second stage, the teacher model uses non-transition emotional train-set from ESD zhou2021seen as prompts to synthesize training samples based on emotion-transition texts generated by GPT-4o (see Appendix C for generation details and examples). We adopt CosyVoice2 du2024cosyvoice2 as the backbone TTS model. The content aligner is composed of five non-causal Transformer layers and two 5×5 convolutional layers with stride 1 and batch normalization ioffe2015batch , following CosyVoice2’s configuration for architectural consistency. In the second stage, the emotion aligner is a lightweight two-layer causal Transformer. The emotional attention bias module includes a linear layer with a hidden dimension of 14 and an MLP output dimension of 7.
@@
In the first stage, the content aligner is trained for 400k steps on 2 NVIDIA 3090 GPUs using Adam kingma2014adam with a learning rate linearly warmed up to 2.5e-4 over the first 10% of steps, then linearly decayed to 0. Each batch contains 90 seconds of speech. In the second stage, the student model is trained for 600k steps on 4 NVIDIA 3090 GPUs. The TTS model is frozen for the first 20k steps to focus on training the emotion aligner. Each batch contains 40 seconds of speech, and Adam is used with a fixed learning rate of 5e-7. Repetition-aware top- sampling chen2024valle2neuralcodec is applied during inference, with and temperature .
@@
To evaluate word-level control over emotion and speaking rate, we construct test sets based on test set of ESD and use outstanding zero-shot TTS models deng2025indextts ; chen2024f5 ; wang2025spark ; du2024cosyvoice2 with multi-round concatenative inference as baselines (see Appendix F for details). We use objective and subjective metrics to assess system performance (see Appendix F.3 for details). For intelligibility, we report WER using Whisper-Large radford2023robust for English and CER using Paraformer gao2022paraformer for Chinese. Speaker similarity (S-SIM) is computed via cosine similarity of WavLM-Large embeddings chen2022wavlm . To evaluate prosody alignment, we use AutoPCP barrault2023seamless . Emotion similarity metrics (Emo2v. and Aro.) are computed using emotion2vec-Large ma2023emotion2vec and a wav2vec-based model baevski2020wav2vec , respectively. We use the variance of DNSMOS-Pro cumlin2024dnsmos (DNSV) to assess the naturalness of emotion transition. Subjective evaluation includes four kinds of Mean Opinion Score (MOS): SMOS (speaker similarity), NMOS (naturalness of emotion transition), EMOS (emotion match), and SPMOS (speed match), each rated on a 5-point scale. Both the mean and 95% confidence intervals of MOS are reported.
@@ 4.2.1 Comparison with Reference Models
We conduct subjective evaluations covering emotional expressiveness (EMOS), speaking rate control (SPMOS), speaker similarity (SMOS), and naturalness of emotion transition (NMOS), with details provided in Appendix F.3. As shown in Table 2, our method, WeSCon, consistently outperforms all baselines. It achieves more expressive and controllable speech while maintaining speaker identity, demonstrating effective word-level control in emotional expression. Additionally, WeSCon delivers more natural-sounding speech with smoother and more accurate speaking rate modulation.
@@
In addition to introducing word-level controllability, we evaluate the performance of our method on the standard zero-shot TTS task using the SEED test set (test-zh) anastassiou2024seedttsfamilyhighqualityversatile . As shown in Table 3, the WeSCon (1st) model yields results identical to CosyVoice2, as the backbone TTS is frozen during this stage. The 2nd-stage model also achieves comparable results. Together with the findings in Table 1, these results demonstrate that our method enables word-level emotion and speaking rate control without significantly degrading the original zero-shot TTS performance of the pretrained model.

### v2

@@ 1 Introduction
Humans possess the ability to regulate emotional expression during speech flexibly Scherer (1995). To simulate this expressive capability, recent advances in text-to-speech synthesis (TTS) have increasingly focused on controllable generation of various aspects of speech, such as timbre, emotion, and speaking rate Xie et al. (2024). Such control is a key objective in the development of human-like and expressive TTS.
Most current TTS models exhibit zero-shot capabilities, enabling them to synthesize speech from text while cloning attributes such as timbre, emotion, and speaking rate from a reference speech sample Kharitonov et al. (2023); Wang et al. (2023); Ju et al. (2024). Despite these advances, as shown in Figure 1, emotional and speaking rate control in current models is typically limited to the utterance level.
This differs significantly from how humans naturally express emotion in speech. Unlike global speaker identity, emotional expression and speaking rate are dynamic and often vary within a single sentence Mozziconacci (2002); Guan et al. (2018). Therefore, word-level control of these factors is essential for achieving more natural and expressive speech synthesis Guo et al. (2022). To address this limitation, some approaches have proposed phoneme-level emotion prediction from target text to guide expressive synthesis Tang et al. (2024); Lei et al. (2021); Du and Yu (2021). While these methods show potential for word-level emotion control, relying solely on text makes it difficult to capture essential acoustic cues such as prosody and intensity, which are vital to emotional expression control Chen and Rudnicky (2022); Chandra et al. (2024); Zhao and Yang (2023). To address this limitation, recent studies such as ELaTE Kanda et al. (2024) and EmoCtrl-TTS Wu et al. (2024) have demonstrated that reference speech with emotional content can support intra-utterance control of time-varying expressive patterns, such as transitions from laughter to crying. These works reflect a growing interest in TTS with word-level control over both emotion and speaking rate, but they also underscore several fundamental challenges. First, word-level expression control requires multiple emotional speech prompts, which introduces the challenge of guiding the model to attend to the appropriate emotion at each word. In addition, current methods for fine-grained expression control rely on large-scale emotional speech datasets with time-aligned emotion transitions. However, such datasets are limited in both scale and accessibility Rathi and Tripathy (2024), making fine-grained control even more difficult to realize in practice. These challenges lead us to ask: Is it possible to achieve effective word-level control of both emotion and speaking rate without relying on speech datasets containing emotion or speed transitions?
@@
We conduct comprehensive experiments to validate the effectiveness of our proposed framework. Results show that our method enables a pretrained zero-shot TTS model to achieve SOTA performance in word-level emotional expression control, while preserving its original zero-shot capabilities. Ablation studies further confirm the contribution of each key design component. Our samples are available at https://wangtianrui.github.io/wescon/.
@@ 2 Related Work
The development of controllable TTS, particularly for emotional expression control, depends heavily on high-quality emotional speech datasets Zhu et al. (2024); Guo et al. (2023); Yang et al. (2025). While public corpora such as ESD Zhou et al. (2021), IEMOCAP Busso et al. (2008), and CREMA-D Cao et al. (2014) are available, they primarily provide utterance-level annotations and lack word-level or time-aligned emotional labels. These datasets are also limited in size and diversity, often consisting of scripted speech and covering a narrow range of emotions and speakers Ma et al. (2024). More importantly, emotional datasets with intra-sentence variation, which are essential for learning word-level control, remain extremely scarce and are typically restricted to private use Kanda et al. (2024). Creating such datasets is expensive, requiring detailed word- or frame-level annotation and subjective emotional labeling Liu et al. (2024). This lack of fine-grained emotional data poses a major challenge for training models capable of word-level expressive TTS.
@@
Most controllable TTS systems support only utterance-level control, where a single label or reference speech governs the entire sentence Anastassiou et al. (2024b); Du et al. (2024a). To achieve word-level control, some methods attempt to predict frame- or phoneme-level emotional indicators from text alone Tan et al. (2024); Im et al. (2022); Luo et al. (2024), but they often fail to capture expressive variability due to the lack of acoustic cues such as intensity and prosody Tang et al. (2024); Lei et al. (2021); Du and Yu (2021). Other approaches, such as ELaTE Kanda et al. (2024) and EmoCtrl-TTS Wu et al. (2024), introduce emotional reference speech to enable intra-utterance control of specific expressive patterns like laughter or crying. While these represent progress, they are typically limited in expressiveness or rely on large-scale emotional datasets that are rarely publicly available. Consequently, achieving general and flexible word-level control over both emotion and speaking rate remains a major challenge.
@@
Self-training has become a promising approach for low-resource speech signal processing, enabling knowledge transfer without fine-grained datasets Zoph et al. (2020); Amini et al. (2025). While it has been applied to tasks like speaker adaptation Khurana et al. (2021), paralinguistic modeling Yang et al. (2024b), and speech translation Pino et al. (2020); Fang et al. (2022), its use for fine-grained emotional control in TTS remains unexplored, especially without detailed expressive labels. To address the scarcity of fine-grained datasets for word-level expressive control, we propose a self-training framework where a teacher model with multi-round inference, transition smoothing, and dynamic speed control generates expressive pseudo-labels. A student model, sharing the teacher’s backbone, is then fine-tuned under its supervision to perform word-level emotion and speaking rate control through a simplified end-to-end inference process, using only a small public dataset without intra-sentence emotion or speed transitions.
@@ 3.2.1 Word-Level Emotion Control
As discussed in Section 2, current TTS models can perform utterance-level emotion and speaker cloning. Building on this, we adopt the high-performance CosyVoice2 Du et al. (2024b) as our backbone (details of the backbone architecture are provided in Appendix A) and propose a multi-round inference strategy, where the model synthesizes multiple segments using different emotional prompts to achieve word-level emotion control. While this approach enables flexible emotional modulation, it often causes unnatural acoustic discontinuities at segment boundaries. To address this, we introduce a transition-smoothing mechanism that improves coherence across inference rounds, as illustrated in Figure 3. Without modifying CosyVoice2, we append a lightweight content aligner, composed of non-causal Transformer Vaswani et al. (2017) and convolutional layers. Trained on ASR data, this module predicts the corresponding text token for each speech token and requires no emotional supervision. During inference, the input text is segmented based on a user-defined emotion plan. At each inference round, the final text and speech tokens from the previous round are appended to the current prompt, forming an explicit tail-to-head linkage. This aligns naturally with CosyVoice2’s continuation-style generation Borsos et al. (2023); Chen et al. (2025), enabling smooth and coherent emotional transitions.
@@ 3.3 Self-Training
In the previous section, we enabled word-level control of emotion and speaking rate by introducing a multi-round inference framework for CosyVoice2 Du et al. (2024b). However, components such as the non-causal content aligner, multi-round inference, and tail-to-head linkage introduce significant inference complexity. To reduce this overhead while preserving controllability, we adopt a self-training strategy. As shown in Figure 4, the enhanced first-stage model serves as a teacher to supervise the original TTS model. The student model, equipped with a dynamic emotional attention bias, learns to achieve word-level emotion and speaking rate control through a simplified end-to-end inference.
@@ 3.3.1 Self-Training with Teacher-Generated Emotion-Transition Speech
Our teacher model achieves word-level control of emotion and speaking rate without modifying the original TTS parameters, relying instead on a complex inference pipeline with dynamic speed control and multi-round generation. To transfer this fine-grained control ability to a simplified end-to-end model, we propose a self-training strategy. Specifically, the 1st-stage teacher model guides the student model to learn word-level controllability. We first use GPT-4o Hurst et al. (2024) to generate emotion-transition text sequences (details are shown in Appendix D), which are paired with public emotional speech samples (without emotion transitions) as prompts. The teacher then synthesizes speech with word-level variation in emotion and speaking rate. These outputs are filtered based on character accuracy and expressive similarity (details are introduced in Appendix E), and the student model is fine-tuned on the filtered supervisions with a small learning rate. This enables word-level emotional expression control during inference without requiring multi-round generation or dynamic concatenation.
@@ 3.3.2 Dynamic Emotional Attention Bias
While the above data formatting preserves CosyVoice2’s generalization by avoiding interference with learned knowledge, it introduces a new challenge: during synthesis, the model may incorrectly attend to emotion-inconsistent prompts. For instance, when generating speech aligned with Emotion I, attention may drift toward prompts labeled with Emotion II, leading to emotional inconsistency and degraded synthesis quality. To address this, we propose a dynamic attention bias mechanism that constrains the model’s focus to emotion-relevant prompt regions based on the predicted emotional trajectory. Concretely, we introduce a causal lightweight Transformer to predict token-level emotion labels for each speech token from historical context. Using the predicted emotion sequence, we introduce a dynamic attention bias mechanism at each Transformer layer. We first concatenate the current text-speech representation with the predicted emotion features and project it through a linear layer. The output is processed in two ways: one path adds a residual and feeds into the next layer, while the other is passed to an MLP Popescu et al. (2009) and softmax to produce a weight vector . The is then used to compute a dynamic attention bias by linearly combining seven predefined attention bias templates (see Appendix F for details). The resulting bias is computed as:
@@ 3.4 Detail Training Setup
We use forced alignment McAuliffe et al. (2017) to generate token-level alignments between transcripts and speech, which serve as supervision for the content aligner. The TTS model remains frozen throughout this stage. Training of the content aligner is conducted without multi-round forwards. Let and denote the input text and speech token sequences, denote the aligned target token sequence, where each label corresponds to one of token classes. Let be the binary label sequence for content boundary detection. The content aligner is jointly trained with a token-level content classification loss and a binary boundary detection loss:
where and denote the last frame indices for text and speech, and is the total number of speech tokens. The learnable parameters correspond to the content aligner, while is the frozen TTS model parameter used during forward propagation. We also apply class weighting during loss computation to reduce the impact of overrepresented silence tokens and address the imbalance in boundary label distribution He and Garcia (2009).
@@ 4.1 Experimental Setup
In the first stage, the content aligner is trained on 200 hours of non-emotional English-Chinese speech from LibriSpeech-100-Clean Panayotov et al. (2015) and AISHELL-1 Bu et al. (2017). In the second stage, the teacher model uses non-transition emotional train-set from ESD Zhou et al. (2021) as prompts to synthesize training samples based on emotion-transition texts generated by GPT-4o (see Appendix D for generation details and examples). We adopt CosyVoice2 Du et al. (2024b) as the backbone TTS model. The content aligner is composed of five non-causal Transformer layers and two 5×5 convolutional layers with stride 1 and batch normalization Ioffe and Szegedy (2015), following CosyVoice2’s configuration for architectural consistency. In the second stage, the emotion aligner is a lightweight two-layer causal Transformer. The emotional attention bias module includes a linear layer with a hidden dimension of 14 and an MLP output dimension of 7.
@@
In the first stage, the content aligner is trained for 400k steps on 2 NVIDIA 3090 GPUs using Adam Kingma (2014) with a learning rate linearly warmed up to 2.5e-4 over the first 10% of steps, then linearly decayed to 0. Each batch contains 90 seconds of speech. In the second stage, the student model is trained for 600k steps on 4 NVIDIA 3090 GPUs. The TTS model is frozen for the first 20k steps to focus on training the emotion aligner. Each batch contains 40 seconds of speech, and Adam is used with a fixed learning rate of 5e-7. Repetition-aware top- sampling Chen et al. (2024a) is applied during inference, with and temperature .
@@
To evaluate word-level control over emotion and speaking rate, we construct test sets based on test set of ESD and use outstanding zero-shot TTS models Deng et al. (2025); Chen et al. (2024b); Wang et al. (2025); Du et al. (2024b) with multi-round concatenative inference as baselines (see Appendix G for details). We use objective and subjective metrics to assess system performance (see Appendix 9 for details). For intelligibility, we report WER using Whisper-Large Radford et al. (2023) for English and CER using Paraformer Gao et al. (2022) for Chinese. Speaker similarity (S-SIM) is computed via cosine similarity of WavLM-Large embeddings Chen et al. (2022). To evaluate prosody alignment, we use AutoPCP Barrault et al. (2023). Emotion similarity metrics (Emo2v. and Aro.) are computed using emotion2vec-Large Ma et al. (2023) and a wav2vec-based model Baevski et al. (2020), respectively. We use the variance of DNSMOS-Pro Cumlin et al. (2024) (DNSV) to assess the naturalness of emotion transition. Subjective evaluation includes four kinds of Mean Opinion Score (MOS): SMOS (speaker similarity), NMOS (naturalness of emotion transition), EMOS (emotion match), and SPMOS (speed match), each rated on a 5-point scale. Both the mean and 95% confidence intervals of MOS are reported.
@@ 4.2.1 Comparison with Reference Models
We conduct subjective evaluations covering emotional expressiveness (EMOS), speaking rate control (SPMOS), speaker similarity (SMOS), and naturalness of emotion transition (NMOS), with details provided in Appendix 9. As shown in Table 2, our method, WeSCon, consistently outperforms all baselines. It achieves more expressive and controllable speech while maintaining speaker identity, demonstrating effective word-level control in emotional expression. Additionally, WeSCon delivers more natural-sounding speech with smoother and more accurate speaking rate modulation.
@@
In addition to introducing word-level controllability, we evaluate the performance of our method on the standard zero-shot TTS task using the SEED test set (test-zh) Anastassiou et al. (2024a). As shown in Table 3, the WeSCon (1st) model yields results identical to CosyVoice2, as the backbone TTS is frozen during this stage. The 2nd-stage model also achieves comparable results. Together with the findings in Table 1, these results demonstrate that our method enables word-level emotion and speaking rate control without significantly degrading the original zero-shot TTS performance of the pretrained model.